OpenStudy (claire.bracken):

For a nonnegative integer n, let r_9(n)stand for the remainder left when n is divided by 9. For example, r_9(25)=7. What is the 22nd entry in an ordered list of all nonnegative integers n that satisfy r_9(5n) =<4? (Note that the first entry in this list is 0.)

OpenStudy (welshfella):

n = 0 r_9(5*0) = 0 n = 1 r_9(5*1) = 5 n = 2 r_9(5*2) = 1 n = 3 r_9(5*3) = 6 so the first 2 are 0, 2

OpenStudy (welshfella):

n = 4 would give 2 n = 5 7 n = 6 3

OpenStudy (claire.bracken):

@welshfella 0, 2, 4, 6... that's a pattern by 2's

OpenStudy (welshfella):

yea looks that way

OpenStudy (claire.bracken):

@welshfella does that mean the answer is 42?

OpenStudy (welshfella):

seems ok when n = 42 r_9(5*42) = 3

OpenStudy (claire.bracken):

@welshfella Alcumus says it's incorrect....

OpenStudy (welshfella):

oh ok lets try some more n = 7 gives 35/9 rem 8 n = 8 gives 40 / 9 rem 4 so n = 8 is not on the list

OpenStudy (welshfella):

n = 9 gives 45/9 rem 0 so n=9 is included in the list so thats why we are wrong. the first five are are 0,2,4,6, 9

OpenStudy (welshfella):

the next one is n = 11 because 55/9 gives rem of 1

OpenStudy (welshfella):

n=13 gives rem 2 n = 15 gives rem 3 so they are going up in odd numbers now

OpenStudy (claire.bracken):

so.. 0, 2, 4, 6, 9, 11, 13, 15...... how do you know when it change? like a pattern to find the 22nd number

OpenStudy (welshfella):

try some more values 16,17 18 ,19 what do we get? this is a tiresome way to answer the question but I cant think of another one at the moment

OpenStudy (welshfella):

16 not included 17 -- 5*17 = 85 85/9 - remainder 4 so 17 is not included

OpenStudy (welshfella):

- maybe we'll get even numbers now.... n=18 5*18 = 90 - rem 0 so n = 18 n=19 - nope n=20 5820 / 9 rem 1 so n = 20

OpenStudy (welshfella):

0,2,4,6,9,11,13,15, 18,20,22,24

OpenStudy (welshfella):

maybe they are in groups of 4 alternate even and odd

OpenStudy (welshfella):

if this is so then the next num ber should be n = 27 n = 25 5*25 = 125 / 9 rem 8 n = 26 5*26 = 130/9 rem 4 n = 27 135 135 /9 rem 0 looks that way after the evens/odds theres a jump of 3

OpenStudy (claire.bracken):

but 26 would also work because the remainder can equal 4... its =<

OpenStudy (welshfella):

no 26 not included

OpenStudy (claire.bracken):

$$r_9(5n)\le 4~?$$

OpenStudy (claire.bracken):

but 26 should work bc remainder 4..^

OpenStudy (welshfella):

the next 4 would be 27 29 31 33 then 36 18 40 42

OpenStudy (claire.bracken):

but the remainder can equal 4 also

OpenStudy (welshfella):

Oh i misread the question I took it as < 4

OpenStudy (claire.bracken):

It's alright :) i think the only change is that it's groups of 5 and a jump of 2?

OpenStudy (claire.bracken):

0, 2, 4, 6, 8, 9, 11, 13, 15, 17, 18, 20, 22, 24, 26, 27, 29, 31, 33, 35, 36, 38, 40, 42, 44 groups of 5 and jump of 1

OpenStudy (welshfella):

yea n = 8 has to be included also n = 17 and n = 26

OpenStudy (claire.bracken):

sothe 22nd would be 38

OpenStudy (claire.bracken):

yup its correct :) answer alcumus gave: The condition $r_9(5n)\le 4$ can also be stated as $``5n\equiv 0,1,2,3,\text{ or }4\pmod 9."$' We can then restate that condition again by multiplying both sides by $2:$ $$10n \equiv 0,2,4,6,\text{ or }8\pmod 9.$$This step is reversible (since $2$ has an inverse modulo $9$). Thus, it neither creates nor removes solutions. Moreover, the left side reduces to $n$ modulo $9,$ giving us the precise solution set $$n \equiv 0,2,4,6,\text{ or }8\pmod 9.$$We wish to determine the $22^{\text{nd}}$ nonnegative integer in this solution set. The first few solutions follow this pattern: $$\begin{array}{c c c c c} 0 & 2 & 4 & 6 & 8 \\ 9 & 11 & 13 & 15 & 17 \\ 18 & 20 & 22 & 24 & 26 \\ 27 & 29 & 31 & 33 & 35 \\ 36 & 38 & \cdots \end{array}$$The $22^{\text{nd}}$ solution is $\boxed{38}.$

OpenStudy (claire.bracken):

thanks so much for all ofyour help they did it basically the same way except mods ;) @welshfella

OpenStudy (welshfella):

thats correct

OpenStudy (welshfella):

yw